OpenStudy (anonymous):

Solve the system by the method of substitution. y = 3x + 8 y = 2x + 7

OpenStudy (anonymous):

So basically you can just subtract the second equation from the first one to get rid of the y. So you get: x+1=0 x=-1 You substitute this into the equation and get y=-3+8 So y=5 Therefore the solution is (-1,5)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

OK, @nincompoop. I guess I forgot the difference, like I always do.